# EMOE: A Framework for Out-of-distribution Uncertainty Based Rejection via Model-Agnostic Expansive Matching of Experts

### v3

@@ Abstract
Expansive Matching of Experts (EMOE) is a novel framework that utilizes support-expanding, extrapolatory pseudo-labeling to improve prediction and uncertainty based rejection on out-of-distribution(OOD) points. EMOE utilizes a diverse set of multiple base experts as pseudo-labelers on the augmented data to improve OOD performance through multiple MLP heads (one per expert) with shared embedding train with a novel per-head matching loss. Unlike prior methods that rely on modality-specific augmentations or assume access to OOD data, EMOE introduces extrapolatory pseudo-labeling on latent-space augmentations, enabling robust OOD generalization with any real-valued vector data. In contrast to prior modality agnostic methods with neural backbones, EMOE is model-agnostic, working effectively with methods from simple tree-based models to complex OOD generalization models. We demonstrate that EMOE achieves superior performance compared to state-of-the-art method on diverse datasets in single-source domain generalization setting.
@@ 1 Introduction
Figure 2: EMOE takes a 2-stage training approach. Stage 1 trains diverse base experts based on different subset of the training data (left). Stage 2 trains Multi-Headed Neural Network with pseudo labels on augmented data provided by the base experts (right).
Contributions  In this work, we propose a new method for single source domain generalization – EMOE, designed to extrapolate effectively and yield reliable uncertainty estimates in high-confidence regions through a novel training scheme using a multi-headed network based on matching augmented (expanded) data (see Fig. 2).
@@
This novel metric measures a model’s ability to predict confidence of true positive examples in its most confident predictions. (4) We show state-of-the-art (SOTA) performance in rejecting predictions based on estimated confidences, evaluated using AU{PR,RO}C-based metrics (see Fig. 1) in a single-source generalization setting (Qiao, Zhao, and Peng 2020). (5) We conduct several ablations to better understand the keys to EMOE’s success; moreover, by ablating the type of base experts, we show EMOE’s broad ability to improve over a model-agnostic set of experts.
@@ 2 Related Work
Self-Training  Self-training uses an earlier model to pseudo-label unlabeled data, which is then added to the training set for subsequent model updates.The concept of pseudo-labeling was initially proposed by (Lee 2013), suggesting a direct approach to retaining instances where the model has high prediction probabilities. Following (Lee 2013), (Zou et al. 2018) proposed selecting a proportion of the most confident unlabeled points instead of using a fixed threshold. Later works combined pseudo-labeling with curriculum learning, dynamically adjusting class-wise thresholds over time to incorporate more informative samples (Cascante-Bonilla et al. 2020; Zhang et al. 2021).Another line of work improves pseudo-labeling robustness by promoting diversity in the labelers. (Ghosh et al. 2021) used model ensembles as teachers, while (Xie et al. 2019) added noise via Dropout (Srivastava et al. 2014) and data augmentation. FixMatch (Sohn et al. 2020) generates pseudo-labels from weakly augmented samples to supervise training on the corresponding strongly augmented samples. EMOE novelly leverages pseudo-labels by assigning each student head to a different expert, encouraging diversity, while aligns the ensemble prediction for consistency via a novel loss.
@@
Ensemble Modeling  Ensembles utilize a diverse set of models jointly for better performance. Early methodologies for ensembles aggregate (bag) predictions from all models (Dietterich 2007; Kussul et al. 2010) or a subset of the models in the ensemble (Jordan and Jacobs 1993; Eigen, Ranzato, and Sutskever 2013). In the OOD setting, prior works addressed this problem by enforcing prediction diversity on OOD data (Pagliardini et al. 2023), ensembling moving average models (Arpit et al. 2022a), training an ensemble of domain specific classifiers (Yao et al. 2023), and training diverse model heads within a single network by maximizing disagreement on unlabeled OOD data (Lee, Yao, and Finn 2022). EMOE adopts a multi-headed architecture that produces an ensemble to improve predictions on OOD data.
@@ 3 Method
Our approach, EMOE, consists of three core components: (1) training a diverse set of base experts; (2) generating extrapolatory samples via latent space augmentation; and (3) training a multi-headed network to match expert predictions on both ID and expanded data. This section describes each step in detail. Throughout, we assume the ‘single-source’ generalization setting (Qiao, Zhao, and Peng 2020), where we observe a single in-distribution (ID) training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and instances are drawn iid $(x_{i},y_{i})\sim\mathcal{P}_{\mathrm{in}}$ without any accompanying environmental/domain/source information nor any labeled/unlabeled OOD instances. For simplicity, we write to the binary classification case, $y_{i}\in\{0,1\}$, but our methodology is easily extendable to other supervised tasks. We design our method to work in general, non-modality specific [^2] We avoid any modality or domain-specific augmentations. (e.g., image, text, audio) settings, i.e., $x_{i}\in\mathbb{R}^{d}$.
@@ 3.1 Base Collection of Experts
EMOE leverages a set of diverse initial experts $\{g_{k}\}_{k=1}^{K}$, s.t. $g_{k}:\mathbb{R}^{d}\rightarrow\{0,1\}$, to guide the training of a secondary model by providing pseudo-labels. There are many mixture of experts (Jordan and Jacobs 1993; Eigen, Ranzato, and Sutskever 2013) and ensembling (Arpit et al. 2022a; Dietterich 2007; Pagliardini et al. 2023; Yao et al. 2023) methods available , EMOE utilizes a collection of strong base-learners by sub-selecting on both instance and feature subspace, encouraging diversity by specializing experts on distinct regions and views of the latent representation space.
@@ 3.2 Expansive Augmentation of Training Data
3.3 EMOE: Expansive Matching of Experts
@@
$\mathcal{L}_{\mathrm{EMOE}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{D})$ (7)
@@
$f_{\mathrm{EMOE}}(x)\equiv\frac{1}{2K}\sum_{j=1}^{K}g_{j}(x)+h_{j}(\phi(x)).$ (11)
@@
We expound on how EMOE may learn better estimates on OOD data through diversity and multi-task learning, and variance reduction and regularization.
@@ 4 Experiments
EMOE base 98.00±0.07 89.48±0.24 99.14±0.02 94.20±0.15 97.79±0.11 68.48±0.30
EMOE 98.45±0.06 89.76±0.26 99.15±0.05 94.42±0.09 98.66±0.10 69.04±0.50
@@
EMOE base 91.21±0.02 82.61±0.05 94.91±0.03 84.13±0.06 88.44±0.05 71.80±0.07
EMOE 91.59±0.03 83.06±0.10 95.38±0.01 84.77±0.02 89.52±0.05 71.41±0.11
We conduct experiments on a varied set of real-world datasets to test the OOD generalizability of EMOE. We considered the single source domain generalization setting (e.g., (Qiao, Zhao, and Peng 2020)), where our model is trained solely on ID data without any (labeled or unlabeled) OOD data during training/validation (e.g., precluding typical semi-supervised approaches), and without any accompanying environmental/domain/source information from ID training instances. Moreover, we note that we avoided utilizing any modality-specific information in EMOE (e.g., we do not utilize any domain specific augmentations) for generality. We utilized XGB Classifiers (Chen and Guestrin 2016) fitted to random subsets of data instances and features as the base collection of experts. For a fair/realistic evaluation, we avoided any hyper-parameter tuning on EMOE and utilized a fixed architecture of a $2$ layer $512$ ELU (Clevert, Unterthiner, and Hochreiter 2015) hidden-unit MLP with $1024$ linear-output heads (please see other hyperparameters in Appx.A.1). For our latent space, we utilize PCA with $128$ components. While OOD generalization is an active field of research (Freiesleben and Grote 2023; Liu et al. 2021), methodology for general (non-modality specific) single source domain generalization is more limited. We provide context to our results through comparisons to a diverse set of existing strong domain generalization methods that approach the problem from various perspectives (and are applicable in the single-source setting).
@@
Semi-Supervised Baselines  We provide further context by comparing to semi-supervised methods for general tabular data, DivDis (Lee, Yao, and Finn 2022) and FixMatch (Sohn et al. 2020). Note, these methods utilize more information than EMOE, and have access to unlabeled OOD data from the test-time distribution. That is, these approaches break from our single-source generalization setting and utilize additional information of the test-time distribution. DivDis is a two-stage framework that first trains diverse model heads using unlabeled target OOD data and then selects the best head with minimal supervision (i.e., using a single label from target OOD data). FixMatch generates filtered pseudo-labels using the model’s predictions on weakly-augmented OOD unlabeled samples. The model is then trained to predict the pseudo-label when provided with a strongly augmented version of the same sample. As DivDis and FixMatch utilize additional information from single-source methods, we separate their results and denote them with an asterisk.
Figure 4: AUPRC at recall << 0.2 for EMOE and competitive baselines on hERG dataset.
Metrics  For prediction thresholding (rejection), we directly utilize the conditional probability $P(Y=1\mid X=x)$ generated by the models. Many real-world applications (e.g. drug virtual screening) utilize only high-confidence predictions. Thus, we paid close attention to high-confidence filtration and reported percent of AUPRC at conservative recall thresholds $\mathrm{AURPC@R}\!\!<\!\!\tau$’ (see Eq.1 for formal definition and Fig.4 for illustration). We also reported AUPRC, AUROC (see Appx. B.1 & B.2 for results on AUROC, which follow a similar trend) We report both base expert (‘EMOE Base’) and EMOE (Eq.11) ensemble performance. Our code will be open-sourced upon publication.
@@
EMOE base 94.44±0.17 98.22±0.08 95.51±0.19 97.84±0.20
EMOE 94.67±0.29 98.87±0.09 96.88±0.25 98.45±0.19
@@
EMOE base 72.19±0.07 84.10±0.01 72.93±0.13 87.43±0.02
EMOE 73.26±0.08 84.60±0.08 73.59±0.15 88.50±0.10
@@ 4.1 Chemical Datasets
To test how well EMOE generalizes to OOD data in chemical domains, we considered seven datasets from ChEMBL (Gaulton et al. 2011), Therapeutics Data Commons (Huang et al. 2021), and DrugOOD (Ji et al. 2023). For all datasets, we represented molecules using extended-connectivity fingerprints (Rogers and Hahn 2010) with radius 2 (ECFP4) and with dimensionality 1024. ECFP4 is a standard method for molecular representation and was chosen for its simplicity in calculation as well as its ability to perform comparably to learned representations, such as those generated by graph neural networks on relevant classification tasks (Zagidullin et al. 2021). Datasets for inhibition of human Ether-à-go-go-Related Gene (hERG), cytotoxicity of human A549 cells (A549_cells), and agonists for Cytochrome P450 2D6 (cyp_2D6) were collected from ChEMBL (Gaulton et al. 2011). For these datasets, binary classification labels were generated with a pChEMBL threshold of 5.0. We also considered another binary classification dataset for Ames mutagenicity (Ames) that was taken from Therapeutics Data Commons (TDC) (Huang et al. 2021). For the ChEMBL and TDC datasets (hERG, A549_cells, cyp_2D6, and Ames), ID and OOD splits were determined based on the Murko scaffold of a molecule, such that OOD data have molecular scaffolds not present in the ID data, mimicking the “scaffold domain” approach utilized in DrugOOD (Ji et al. 2023).
@@
Figure 5: Predicted probabilities from EMOE network and base experts. We highlight example instances where the base experts initially makes incorrect predictions but are corrected when we average the predicted probabilities from EMOE network and EMOE base.
Results  We assessed models on ChEMBL, TDC, and DrugOOD datasets; our results are shown in Tab. 2 and Tab. 1. Across the chemical datasets, EMOE consistently outperforms both supervised and semi-supervised baselines and consistently shows gains over its base experts. Note further, that EMOE’s performance gain is especially impactful at conservative recalls (AUPRC@R<<0.2, e.g., see Fig. 4), which indicates the EMOE outperform other models in virtual screening tasks to filter drug candidates (where false positives would lead to wasted resources). Please see discussion below ($\S$ 4.4), for further analysis on the improved performance of EMOE w.r.t. increased confidence (Fig. 5) and expert diversity (Fig. 6).
@@ 4.2 Other Real World Datasets
EMOE base 97.39±0.06 90.07±0.32 58.30±0.09 78.62±1.57
EMOE 97.92±0.20 91.72±0.62 67.57±0.12 76.95±1.45
@@
EMOE base 86.39±0.19 83.80±0.07 62.83±0.03 64.01±0.94
EMOE 86.70±0.28 84.02±0.09 63.62±0.08 62.21±0.52
Next, we further evaluate our method in non-chemical domains across a diverse range of real-world OOD scenarios using the Tableshift datasets (Gardner, Popovic, and Schmidt 2023) (and the PACS image dataset (Li et al. 2017), see  Mat. B.5). We selected a diverse collection of Tableshift datasets, based on unrestricted availability and in/out-of-domain performance discrepancy, coverings areas including: finance, education, and healthcare. Each dataset has an associated real-world shift and a related prediction target (see Gardner, Popovic, and Schmidt (2023) for further details). Results on the Tableshift are shown in Tab. 8. As before, we consider the same single-source domain generalization setting. We can see that even over diverse applications, our EMOE method is able to perform well and is often outperforming our strong competing baselines. Moreover, eventhough DivDis and FixMatch use additional unlabeled OOD data, EMOE outperforms both on the majority of datasets, highlighting EMOE’s ability for generalization on completely unforeseen OOD instances (without any knowledge of test time distribution).

### v4

# EXPLOR: Extrapolatory Pseudo-Label Matching for Out-of-distribution Uncertainty Based Rejection

@@ Abstract
EXPLOR is a novel framework that utilizes support-expanding, extrapolatory pseudo-labeling to improve prediction and uncertainty-based rejection on out-of-distribution (OOD) points. EXPLOR utilizes a diverse set of base models as pseudo-labelers on the expansive augmented data to improve OOD performance through multiple MLP heads (one per base model) with shared embedding trained with a novel per-head matching loss. Unlike prior methods that rely on modality-specific augmentations or assume access to OOD data, EXPLOR introduces extrapolatory pseudo-labeling on latent-space augmentations, enabling robust OOD generalization with any real-valued vector data. In contrast to prior modality-agnostic methods with neural backbones, EXPLOR is model-agnostic, working effectively with methods from simple tree-based models to complex OOD generalization models. We demonstrate that EXPLOR achieves superior performance compared to state-of-the-art methods on diverse datasets in single-source domain generalization settings.
@@ 1 Introduction
Figure 2: EXPLOR takes a 2-stage training approach. Stage 1 trains diverse base experts based on different subset of the training data (left). Stage 2 trains Multi-Headed Neural Network with pseudo labels on augmented data provided by the base experts (right).
Contributions  In this work, we propose a new method for single source domain generalization – EXPLOR, designed to extrapolate effectively and yield reliable uncertainty estimates in high-confidence regions through a novel training scheme using a multi-headed network based on matching augmented (expanded) data (see Fig. 2).
@@
This novel metric measures a model’s ability to predict confidence of true positive examples in its most confident predictions. (4) We show state-of-the-art (SOTA) performance in rejecting predictions based on estimated confidences, evaluated using AU{PR,RO}C-based metrics (see Fig. 1) in a single-source generalization setting (Qiao, Zhao, and Peng 2020). (5) We conduct several ablations to better understand the keys to EXPLOR’s success; moreover, by ablating the type of base experts, we show EXPLOR’s broad ability to improve over a model-agnostic set of experts.
@@ 2 Related Work
Self-Training  Self-training uses an earlier model to pseudo-label unlabeled data, which is then added to the training set for subsequent model updates.The concept of pseudo-labeling was initially proposed by (Lee 2013), suggesting a direct approach to retaining instances where the model has high prediction probabilities. Following (Lee 2013), (Zou et al. 2018) proposed selecting a proportion of the most confident unlabeled points instead of using a fixed threshold. Later works combined pseudo-labeling with curriculum learning, dynamically adjusting class-wise thresholds over time to incorporate more informative samples (Cascante-Bonilla et al. 2020; Zhang et al. 2021).Another line of work improves pseudo-labeling robustness by promoting diversity in the labelers. (Ghosh et al. 2021) used model ensembles as teachers, while (Xie et al. 2019) added noise via Dropout (Srivastava et al. 2014) and data augmentation. FixMatch (Sohn et al. 2020) generates pseudo-labels from weakly augmented samples to supervise training on the corresponding strongly augmented samples. EXPLOR novelly leverages pseudo-labels by assigning each student head to a different expert, encouraging diversity, while aligns the ensemble prediction for consistency via a novel loss.
@@
Ensemble Modeling  Ensembles utilize a diverse set of models jointly for better performance. Early methodologies for ensembles aggregate (bag) predictions from all models (Dietterich 2007; Kussul et al. 2010) or a subset of the models in the ensemble (Jordan and Jacobs 1993; Eigen, Ranzato, and Sutskever 2013). In the OOD setting, prior works addressed this problem by enforcing prediction diversity on OOD data (Pagliardini et al. 2023), ensembling moving average models (Arpit et al. 2022a), training an ensemble of domain specific classifiers (Yao et al. 2023), and training diverse model heads within a single network by maximizing disagreement on unlabeled OOD data (Lee, Yao, and Finn 2022). EXPLOR adopts a multi-headed architecture that produces an ensemble to improve predictions on OOD data.
@@ 3 Method
Our approach, EXPLOR, consists of three core components: (1) training a diverse set of base experts; (2) generating extrapolatory samples via latent space augmentation; and (3) training a multi-headed network to match expert predictions on both ID and expanded data. This section describes each step in detail. Throughout, we assume the ‘single-source’ generalization setting (Qiao, Zhao, and Peng 2020), where we observe a single in-distribution (ID) training dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, and instances are drawn iid $(x_{i},y_{i})\sim\mathcal{P}_{\mathrm{in}}$ without any accompanying environmental/domain/source information nor any labeled/unlabeled OOD instances. For simplicity, we write to the binary classification case, $y_{i}\in\{0,1\}$, but our methodology is easily extendable to other supervised tasks. We design our method to work in general, non-modality specific [^2] We avoid any modality or domain-specific augmentations. (e.g., image, text, audio) settings, i.e., $x_{i}\in\mathbb{R}^{d}$.
@@ 3.1 Base Collection of Experts
EXPLOR leverages a set of diverse initial experts $\{g_{k}\}_{k=1}^{K}$, s.t. $g_{k}:\mathbb{R}^{d}\rightarrow\{0,1\}$, to guide the training of a secondary model by providing pseudo-labels. There are many mixture of experts (Jordan and Jacobs 1993; Eigen, Ranzato, and Sutskever 2013) and ensembling (Arpit et al. 2022a; Dietterich 2007; Pagliardini et al. 2023; Yao et al. 2023) methods available , EXPLOR utilizes a collection of strong base-learners by sub-selecting on both instance and feature subspace, encouraging diversity by specializing experts on distinct regions and views of the latent representation space.
@@ 3.2 Expansive Augmentation of Training Data
3.3 EXPLOR: Expansive Matching of Experts
@@
$\mathcal{L}_{\mathrm{EXPLOR}}(\phi,\{h_{j}\}_{j=1}^{K},\{g_{j}\}_{j=1}^{K};\mathcal{D})$ (7)
@@
$f_{\mathrm{EXPLOR}}(x)\equiv\frac{1}{2K}\sum_{j=1}^{K}g_{j}(x)+h_{j}(\phi(x)).$ (11)
@@
We expound on how EXPLOR may learn better estimates on OOD data through diversity and multi-task learning, and variance reduction and regularization.
@@ 4 Experiments
EXPLOR base 98.00±0.07 89.48±0.24 99.14±0.02 94.20±0.15 97.79±0.11 68.48±0.30
EXPLOR 98.45±0.06 89.76±0.26 99.15±0.05 94.42±0.09 98.66±0.10 69.04±0.50
@@
EXPLOR base 91.21±0.02 82.61±0.05 94.91±0.03 84.13±0.06 88.44±0.05 71.80±0.07
EXPLOR 91.59±0.03 83.06±0.10 95.38±0.01 84.77±0.02 89.52±0.05 71.41±0.11
We conduct experiments on a varied set of real-world datasets to test the OOD generalizability of EXPLOR. We considered the single source domain generalization setting (e.g., (Qiao, Zhao, and Peng 2020)), where our model is trained solely on ID data without any (labeled or unlabeled) OOD data during training/validation (e.g., precluding typical semi-supervised approaches), and without any accompanying environmental/domain/source information from ID training instances. Moreover, we note that we avoided utilizing any modality-specific information in EXPLOR (e.g., we do not utilize any domain specific augmentations) for generality. We utilized XGB Classifiers (Chen and Guestrin 2016) fitted to random subsets of data instances and features as the base collection of experts. For a fair/realistic evaluation, we avoided any hyper-parameter tuning on EXPLOR and utilized a fixed architecture of a $2$ layer $512$ ELU (Clevert, Unterthiner, and Hochreiter 2015) hidden-unit MLP with $1024$ linear-output heads (please see other hyperparameters in Appx.LABEL:EXPLOR_hp). For our latent space, we utilize PCA with $128$ components. While OOD generalization is an active field of research (Freiesleben and Grote 2023; Liu et al. 2021), methodology for general (non-modality specific) single source domain generalization is more limited. We provide context to our results through comparisons to a diverse set of existing strong domain generalization methods that approach the problem from various perspectives (and are applicable in the single-source setting).
@@
Semi-Supervised Baselines  We provide further context by comparing to semi-supervised methods for general tabular data, DivDis (Lee, Yao, and Finn 2022) and FixMatch (Sohn et al. 2020). Note, these methods utilize more information than EXPLOR, and have access to unlabeled OOD data from the test-time distribution. That is, these approaches break from our single-source generalization setting and utilize additional information of the test-time distribution. DivDis is a two-stage framework that first trains diverse model heads using unlabeled target OOD data and then selects the best head with minimal supervision (i.e., using a single label from target OOD data). FixMatch generates filtered pseudo-labels using the model’s predictions on weakly-augmented OOD unlabeled samples. The model is then trained to predict the pseudo-label when provided with a strongly augmented version of the same sample. As DivDis and FixMatch utilize additional information from single-source methods, we separate their results and denote them with an asterisk.
Figure 4: AUPRC at recall << 0.2 for EXPLOR and competitive baselines on hERG dataset.
Metrics  For prediction thresholding (rejection), we directly utilize the conditional probability $P(Y=1\mid X=x)$ generated by the models. Many real-world applications (e.g. drug virtual screening) utilize only high-confidence predictions. Thus, we paid close attention to high-confidence filtration and reported percent of AUPRC at conservative recall thresholds $\mathrm{AURPC@R}\!\!<\!\!\tau$’ (see Eq.1 for formal definition and Fig.4 for illustration). We also reported AUPRC, AUROC (see Appx. B.1 & B.2 for results on AUROC, which follow a similar trend) We report both base expert (‘EXPLOR Base’) and EXPLOR (Eq.11) ensemble performance. Our code will be open-sourced upon publication.
@@
EXPLOR base 94.44±0.17 98.22±0.08 95.51±0.19 97.84±0.20
EXPLOR 94.67±0.29 98.87±0.09 96.88±0.25 98.45±0.19
@@
EXPLOR base 72.19±0.07 84.10±0.01 72.93±0.13 87.43±0.02
EXPLOR 73.26±0.08 84.60±0.08 73.59±0.15 88.50±0.10
@@ 4.1 Chemical Datasets
To test how well EXPLOR generalizes to OOD data in chemical domains, we considered seven datasets from ChEMBL (Gaulton et al. 2011), Therapeutics Data Commons (Huang et al. 2021), and DrugOOD (Ji et al. 2023). For all datasets, we represented molecules using extended-connectivity fingerprints (Rogers and Hahn 2010) with radius 2 (ECFP4) and with dimensionality 1024. ECFP4 is a standard method for molecular representation and was chosen for its simplicity in calculation as well as its ability to perform comparably to learned representations, such as those generated by graph neural networks on relevant classification tasks (Zagidullin et al. 2021). Datasets for inhibition of human Ether-à-go-go-Related Gene (hERG), cytotoxicity of human A549 cells (A549_cells), and agonists for Cytochrome P450 2D6 (cyp_2D6) were collected from ChEMBL (Gaulton et al. 2011). For these datasets, binary classification labels were generated with a pChEMBL threshold of 5.0. We also considered another binary classification dataset for Ames mutagenicity (Ames) that was taken from Therapeutics Data Commons (TDC) (Huang et al. 2021). For the ChEMBL and TDC datasets (hERG, A549_cells, cyp_2D6, and Ames), ID and OOD splits were determined based on the Murko scaffold of a molecule, such that OOD data have molecular scaffolds not present in the ID data, mimicking the “scaffold domain” approach utilized in DrugOOD (Ji et al. 2023).
@@
Figure 5: Predicted probabilities from EXPLOR network and base experts. We highlight example instances where the base experts initially makes incorrect predictions but are corrected when we average the predicted probabilities from EXPLOR network and EXPLOR base.
Results  We assessed models on ChEMBL, TDC, and DrugOOD datasets; our results are shown in Tab. 2 and Tab. 1. Across the chemical datasets, EXPLOR consistently outperforms both supervised and semi-supervised baselines and consistently shows gains over its base experts. Note further, that EXPLOR’s performance gain is especially impactful at conservative recalls (AUPRC@R<<0.2, e.g., see Fig. 4), which indicates the EXPLOR outperform other models in virtual screening tasks to filter drug candidates (where false positives would lead to wasted resources). Please see discussion below ($\S$ 4.4), for further analysis on the improved performance of EXPLOR w.r.t. increased confidence (Fig. 5) and expert diversity (Fig. 6).
@@ 4.2 Other Real World Datasets
EXPLOR base 97.39±0.06 90.07±0.32 58.30±0.09 78.62±1.57
EXPLOR 97.92±0.20 91.72±0.62 67.57±0.12 76.95±1.45
@@
EXPLOR base 86.39±0.19 83.80±0.07 62.83±0.03 64.01±0.94
EXPLOR 86.70±0.28 84.02±0.09 63.62±0.08 62.21±0.52
Next, we further evaluate our method in non-chemical domains across a diverse range of real-world OOD scenarios using the Tableshift datasets (Gardner, Popovic, and Schmidt 2023) (and the PACS image dataset (Li et al. 2017), see  Mat. B.5). We selected a diverse collection of Tableshift datasets, based on unrestricted availability and in/out-of-domain performance discrepancy, coverings areas including: finance, education, and healthcare. Each dataset has an associated real-world shift and a related prediction target (see Gardner, Popovic, and Schmidt (2023) for further details). Results on the Tableshift are shown in Tab. 8. As before, we consider the same single-source domain generalization setting. We can see that even over diverse applications, our EXPLOR method is able to perform well and is often outperforming our strong competing baselines. Moreover, eventhough DivDis and FixMatch use additional unlabeled OOD data, EXPLOR outperforms both on the majority of datasets, highlighting EXPLOR’s ability for generalization on completely unforeseen OOD instances (without any knowledge of test time distribution).